OpenStudy (anonymous):

please help!!!!!!!!!!! fan and medal Solve the equation cos( θ+60◦)=2sinθ

OpenStudy (perl):

solve for theta?

OpenStudy (perl):

$$ \Large { \cos(\theta+60~^o)=2\sinθ \\\iff \\ \cos(\theta) \cos (60~^o)- \sin(\theta) \sin (60~^o) =2\sinθ } $$